Jan is a 13-year old exchange student from Argentina. Jan's father is 44 and his mother is 50. Jan's brother Dan is 11 years old. In how many years will Jan be half his mother's age? (Don't forget Jan's mother ages too)
Answer by macston(5194)

.
J=Jan's age; M=mother's age
.
x=years until mother is twice Jan's age.
.
M+x=2(J+x)
50+x=2(13+x)
50+x=26+2x
24=x
ANSWER: Jan will be one half his mothers age in 24 years.